We are given 8 integer numbers. Count the number of non-empty sets which contain at
most 6 of these numbers (and no other numbers).
Answer #1

number of non-empty sets which contain at most 6 of these numbers

=P(contain 1 number)+P(contain 2 number)+P(contain 3 number)+P(contain 4 number)+P(contain 5 number)+P(contain 6 number)=(8C1)+(8C2)+(8C3)+(8C4)+(8C5)+(8C6) =8+28+56+70+56+28 =246